# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v1

@@ H.5 Impact of sampling optimization using discrete flow matching
The results for the synthetic datasets in Figure 4 illustrate that the sampling hyperparameters meaningfully influence the generative performance. We observe that different values can lead to notable variations in both V.U.N. and the structural ratio. However, a higher V.U.N. does not necessarily correspond to a lower ratio. While certain configurations tend to perform well across both RRWP and MagLap positional encodings, the improvements are generally modest, suggesting that optimal settings may vary slightly depending on the dataset and positional encoding strategy used.
@@ I.1 Graph iterative refinement methods
Within the realm of undirected graph generation, discrete state-space iterative refinement approaches vignac2023digress ; xu2024disco ; siraudin2024cometh ; qin2025defog have emerged as a powerful framework for capturing the intricate dependencies that govern target graph distributions. These methods achieve state-of-the-art performance by naturally aligning with the structure of graphs: they operate directly in discrete state spaces, matching the inherent discreteness of adjacency matrices, and respect node permutation equivariances due to their one-shot prediction formulation. These characteristics make iterative refinement approaches particularly well-suited for the task of directed graph generation.
@@
which predicts categorical distributions over nodes and edges given a noised graph $G_{t}$. In practice, $\mathbf{p}^{\theta}_{1|t}(\cdot|G_{t})$ is parameterized using a graph transformer such as dwivedi2021graphtransformer , which allows for expressive modeling of complex graph structures. The network is trained using a cross-entropy loss applied independently to each node and each edge:
@@ I.2 Discrete flow matching for graph generation
Among graph iterative refinement methods, Discrete Flow Matching (DFM) campbell2024generative ; gat2024discrete has recently emerged as a particularly powerful framework. By decoupling the training and sampling processes, DFM enables a broader and more flexible design space, which has been shown to improve generative performance. This framework has proven particularly effective for undirected graph generation qin2025defog .
@@
+\;+ $\sum_{i\neq j}\delta(G_{t}^{\backslash(i,j)},G_{t+\Delta t}^{\backslash(i,j)})\,\mathbb{E}_{p_{1\mid t}^{\theta,(i,j)}(e_{1}^{(i,j)}\mid G_{t})}\left[R_{t}^{(i,j)}(e_{t}^{(i,j)},e_{t+\Delta t}^{(i,j)}\mid e_{1}^{(i,j)})\right],$ (46)
@@ I.3 Discrete diffusion for graph generation
Among discrete diffusion-based models for undirected graphs, DiGress vignac2023digress , grounded in the structured discrete diffusion framework austin2021structured , has been particularly influential. This approach mostly differs from the DFM-based formulation described in Sec. I.2, as it operates in discrete time, with the denoising process modeled as a Discrete-Time Markov Chain (DTMC), in contrast to the continuous-time formulation used in DFM.
The forward, or noising, process is modeled as a Markov noise process $q$, which generates a sequence of progressively noised graphs $G_{t}$, for $t=1,\ldots,T$. At each timestep, node and edge tensors are perturbed using categorical transition matrices $[\mathbf{Q}^{X}_{t}]^{(i,j)}=q(x_{t}=j\mid x_{t-1}=i)$ and $[\mathbf{Q}^{E}_{t}]^{(i,j)}=q(e_{t}=j\mid e_{t-1}=i)$, respectively. This process induces structural changes such as edge addition or deletion, and edits to node and edge categories. The transition probabilities can be summarized as:
@@
where $\alpha^{t}$ transitions from 1 to 0 with $t$ according to the popular cosine scheduling nichol2021improved . The vectors $\mathbf{1}_{X}\in\{1\}^{X}$ and $\mathbf{1}_{E}\in\{1\}^{E+1}$ are filled with ones, and $\mathbf{m}_{X}\in\Delta^{X}$ and $\mathbf{m}_{E}\in\Delta^{E+1}$ are vectors filled with the marginal node and edge distributions, respectively [^2] We denote the probability simplex of the state-space of cardinality $Z$ by $\Delta^{Z}$.
This noising process can be rewritten using the analogous notation to the one used for DFM in Eq. 49 as:
@@
and similarly for edges. To compute the missing posterior term in Eq. 51, $p_{t^{\prime}+\Delta t^{\prime}|1,t^{\prime}}(x^{(n)}_{t^{\prime}+\Delta t^{\prime}}|x^{(n)}_{1},G_{t^{\prime}})$, we equate it to the posterior term of the forward process:
@@
Appendix K Impact statement and limitations
The objective of this work is to advance methodologies for the generation of directed graphs by introducing architectural mechanisms that explicitly model directionality and asymmetric relationships. Directed graphs are central to a wide range of applications, including causal inference, traffic modeling, and biological network analysis. Accordingly, improvements in directed graph generation have the potential to support progress in scientific research, decision-making systems, infrastructure design, or causal discovery.
The proposed framework enhances the expressiveness of generative models to directed graphs, and, at the moment, we do not identify any immediate societal risks associated with its deployment in its current form. Moreover, although the method is capable of generating structured and semantically meaningful graphs, the scale and complexity of these outputs remain limited, which may difficult direct applicability in high-impact domains such as real-time clinical diagnostics or large-scale policy modeling at this stage.

### v2

@@ H.8 Impact of sampling optimization using discrete flow matching
The results for the synthetic datasets in Figure 4 illustrate that the sampling hyperparameters influence the generative performance. We observe that different values can lead to variations in both V.U.N. and the structural ratio. However, a higher V.U.N. does not necessarily correspond to a lower ratio. While certain configurations tend to perform well across both RRWP and MagLap positional encodings, the improvements are generally modest, suggesting that optimal settings may vary slightly depending on the dataset and positional encoding strategy used.
@@ I.1 Graph iterative refinement methods
Within the realm of undirected graph generation, discrete state-space iterative refinement approaches (Vignac et al., 2023a; Xu et al., 2024; Siraudin et al., 2024; Qin et al., 2025a) have emerged as a powerful framework for capturing the intricate dependencies that govern target graph distributions. These methods achieve state-of-the-art performance by naturally aligning with the structure of graphs: they operate directly in discrete state spaces, matching the inherent discreteness of adjacency matrices, and respect node permutation equivariances due to their one-shot prediction formulation. These characteristics make iterative refinement approaches particularly well-suited for the task of directed graph generation.
@@
which predicts categorical distributions over nodes and edges given a noised graph $G_{t}$. In practice, $\mathbf{p}^{\theta}_{1|t}(\cdot|G_{t})$ is parameterized using a graph transformer such as the one from Dwivedi & Bresson (2021), which allows for expressive modeling of complex graph structures. The network is trained using a cross-entropy loss applied independently to each node and each edge:
@@ I.2 Discrete flow matching for graph generation
Among graph iterative refinement methods, Discrete Flow Matching (DFM) (Campbell et al., 2024; Gat et al., 2024) has recently emerged as a particularly powerful framework. By decoupling the training and sampling processes, DFM enables a broader and more flexible design space, which has been shown to improve generative performance. This framework has proven particularly effective for undirected graph generation (Qin et al., 2025a).
@@
$+\;$ $\sum_{i\neq j}\delta(G_{t}^{\backslash(i,j)},G_{t+\Delta t}^{\backslash(i,j)})\,\mathbb{E}_{p_{1\mid t}^{\theta,(i,j)}(e_{1}^{(i,j)}\mid G_{t})}\left[R_{t}^{(i,j)}(e_{t}^{(i,j)},e_{t+\Delta t}^{(i,j)}\mid e_{1}^{(i,j)})\right],$ (43)
@@ I.3 Discrete diffusion for graph generation
Among discrete diffusion-based models for undirected graphs, DiGress (Vignac et al., 2023a), grounded in the structured discrete diffusion framework (Austin et al., 2021), has been particularly influential. This approach mostly differs from the DFM-based formulation described in Section I.2, as it operates in discrete time, with the denoising process modeled as a Discrete-Time Markov Chain (DTMC), in contrast to the continuous-time formulation used in DFM.
The forward, or noising, process is modeled as a Markov noise process $q$, which generates a sequence of progressively noised graphs $G_{t}$, for $t=1,\ldots,T$. At each timestep, node and edge tensors are perturbed using categorical transition matrices $\mathbf{Q}^{X}_{t}]^{(i,j)}=q(x_{t}=j\mid x_{t-1}=i)$ and $\mathbf{Q}^{E}_{t}]^{(i,j)}=q(e_{t}=j\mid e_{t-1}=i)$, respectively. This process induces structural changes such as edge addition or deletion, and edits to node and edge categories. The transition probabilities can be summarized as:
@@
where $\alpha^{t}$ transitions from 1 to 0 with $t$ according to the popular cosine scheduling (Nichol & Dhariwal, 2021). The vectors $\mathbf{1}_{X}\in\{1\}^{X}$ and $\mathbf{1}_{E}\in\{1\}^{E+1}$ are filled with ones, and $\mathbf{m}_{X}\in\Delta^{X}$ and $\mathbf{m}_{E}\in\Delta^{E+1}$ are vectors filled with the marginal node and edge distributions, respectively [^2] We denote the probability simplex of the state-space of cardinality $Z$ by $\Delta^{Z}$.
This noising process can be rewritten using the analogous notation to the one used for DFM in Equation 46 as:
@@
and similarly for edges. To compute the missing posterior term in Equation 48, $p_{t^{\prime}+\Delta t^{\prime}|1,t^{\prime}}(x^{(n)}_{t^{\prime}+\Delta t^{\prime}}|x^{(n)}_{1},G_{t^{\prime}})$, we equate it to the posterior term of the forward process:
@@
Appendix K Limitations and future work
Although Directo shows strong performance in the generation of directed graphs, several avenues for future work could further enhance its scalability, controllability, and applicability.
Scalability to Large Graphs and Datasets:
Directed graphs present a combinatorial explosion in the space of possible edges due to asymmetry, posing challenges in both training and sampling efficiency. Future work could explore sparsity-aware attention mechanisms (Qin et al., 2025b) to reduce quadratic memory and computational complexity. Hierarchical or latent diffusion approaches (Bergmeister et al., 2024; Jang et al., 2024; Yang et al., 2024) may allow the model to efficiently generate large-scale graphs by operating at multiple levels of granularity or in a lower-dimensional latent space. Investigating adaptive sampling schedules and dataset subsampling strategies could also help manage computational costs without sacrificing model quality.
Conditional Graph Generation:
Many real-world applications require generating graphs conditioned on specific properties or context, such as node features, initial subgraphs, or global graph characteristics. DIRECTO already supports classifier-free conditional generation, which allows flexible conditioning without requiring additional networks. However, this approach may not be optimal for all tasks, and exploring alternative conditional strategies (such as learned conditional embeddings or auxiliary networks) could further improve performance. Such extensions would enable targeted graph design, e.g., generating traffic networks with predefined entry/exit nodes, molecular graphs with given scaffolds, or DAGs adhering to specific causal structures
Explicit Structural Constraints:
While Directo can implicitly learn structural constraints such as acyclicity, certain domains require strict adherence to properties like DAG structure, planarity, or node degree limits. Integrating ideas from methods such as ConStrucy (Madeira et al., 2024) or PRODIGY (Sharma et al., 2024) could enforce these constraints explicitly, enhancing control over generated graphs. Future research could also explore soft vs. hard constraint integration, allowing a trade-off between flexibility and domain-specific validity.
Integration with Downstream Tasks:
Beyond purely generative evaluation, integrating Directo with downstream tasks such as causal inference, traffic simulation, or molecular optimization—could highlight practical benefits and guide model improvements. Extending the metrics in the benchmark to further reflect task-specific objectives could be relevant to quantify real-world impact.
Interpretability and Controllability:
Understanding which components (e.g., dual attention vs. directional positional encodings) drive specific structural properties in the generated graphs is an open research question. Future work could investigate interpretable latent representations and mechanisms to control generation in a predictable manner.